which equation of the line has a y-intercept of 6 and is a parallel to y=-1/2+2?
Answer by uma(370)

I believe the equation of the line to be y = -1/2 x + 2
Slope of this line = -1/2 [In y = mx + c, m is the slope]
So slope of any line parallel to this line = -1/2 [Parallel lines have same slope]
Y intercept is given to be 6.
The required equation is y = mx + c
==> y = -1/2 x + 6
Or x + 2y = 5 [on simplification]
Good Luck !!!